# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 2.2 Runtime and branching framework
We say branch-seq $B^{\prime}$ dominates $B$ if $\operatorname{val}_{a,b}(B^{\prime})\leq\operatorname{val}_{a,b}(B)$ for all $a,b\geq 0$, and $G$ has a branch-seq $B$ available if we can efficiently find a branching rule which has some branch-seq $B^{\prime}$ dominating $B$. In general, a “balanced” branch-seq dominates an “imbalanced” one; formally, if $\Delta k_{1}\leq\Delta k_{2}$ and $\Delta\mu_{1}\leq\Delta\mu_{2}$, then branch-seq $[(\Delta\mu_{1},\Delta k_{1}),(\Delta\mu_{2},\Delta k_{2})]$ dominates $[(\Delta\mu_{1}-f,\Delta k_{1}-g),(\Delta\mu_{1}+f,\Delta k_{2}+g)]$ for any $f,g\geq 0$.
We use the following important compositional property throughout: if we have a branching rule $B$ generating $\ell$ subproblems with drops $(\Delta\mu_{i},\Delta k_{i})$, and we then apply a branching rule with branch-seq $B^{\prime}_{i}$ to each subproblem $i=1,\dots,\ell$, then, overall, we get
$\operatorname{val}_{a,b}(B)=\sum_{i=1}^{\ell}e^{-a\Delta\mu_{i}-b\Delta k_{i}}\operatorname{val}_{a,b}(B_{i}^{\prime}).$
Lemma 10.
3 Preprocessing rules and reductions in $k$
We define $S(G)$ to be the largest decrease in $k$ obtainable from applying rules (P1) – (P3) exhaustively via some efficiently-computable sequence of operations. [^3] It is not clear how to calculate the absolute largest decrease in $k$ via preprocessing rules, since applying some rules may prevent opportunities for other rules. We assume that we have fixed some polynomial-time computable sequences of potential preprocessing rules in order to reduce $k$ as small as much as possible. At various points in our algorithm, we will describe simplifications available for intermediate graphs. We always assume that our preprocessing rules are chosen to find such simplifications. As a point of notation, we define $R(X)=S(G-X)$ for a vertex set $X$. For vertices $x_{1},\dots,x_{\ell}$ and vertex sets $X_{1},\dots,X_{r}$, we write $R(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for the more cumbersome $R(\{x_{1},\dots,x_{\ell}\}\cup X_{1}\cup\dots\cup X_{r})=S(G-\{x_{1},\dots,x_{\ell}\}-(X_{1}\cup\dots\cup X_{r}))$. Note that we may have $\ell=0$ or $r=0$, e.g. $R(x_{1},\dots,x_{\ell})$ is shorthand for $R(\{x_{1},\dots,x_{\ell}\})$.
Observation 11.
Suppose $x,y$ are non-adjacent subquadratic vertices. Then $S(G)\geq 2$ if any of the following three conditions hold: (i) $\operatorname{codeg}(x,y)=0$ or (ii) $\deg(x)+\deg(y)\geq 3$ or (iii) $\operatorname{minsurp}(G)\geq 0$.
@@ 3 Preprocessing rules and reductions in $k$
If $G$ has 2-vertices $x_{1},\dots,x_{\ell}$ which all have pairwise distance at least $3$, then $S(G)\geq\ell$.
@@
Lemma 13.
If $\operatorname{minsurp}(G)\geq 0$ and indset $I$ has $\operatorname{surp}_{G}(I)\leq 1$, then $S(G)\geq|I|$.
@@
Let $r=|I|$. If $\operatorname{surp}_{G}(I)=0$, or $I$ is a critical-set, we can apply (P1) or (P2) to $I$. So suppose that $\operatorname{surp}_{G}(I)=1$ and $\operatorname{surp}_{G}(X)=0$ for a non-empty subset $X\subsetneq I$; among all such subsets $X$, choose one of maximum size, and let $s=|X|$. We apply (P1) to $X$, reducing $k$ by $s$ and obtaining a graph $G^{\prime}=G-N[X]$. Now consider the indset $J=I\setminus X$ in $G^{\prime}$, where $|J|=r-s$. We have $|N_{G^{\prime}}(J)|=|N_{G}(I)\setminus N_{G}(X)|=r+1-s$ and so $\operatorname{surp}_{G^{\prime}}(J)=1$.
We claim that $J$ is a critical-set in $G^{\prime}$. For, if some non-empty subset $X^{\prime}\subseteq J$ has $|N_{G^{\prime}}(X^{\prime})|\leq|X^{\prime}|$, then indset $I^{\prime}=X\cup X^{\prime}\subseteq I$ would have $|N_{G}(I^{\prime})|\leq|N_{G}(X)|+|N_{G^{\prime}}(X^{\prime})|\leq|X|+|X^{\prime}|=|I^{\prime}|$. Hence $\operatorname{surp}_{G}(I^{\prime})\leq 0$, contradicting maximality of $X$. So, we can apply (P2) to $J$ in $G^{\prime}$, getting a further drop of $r-s$. Overall, we get a net drop of $s+(r-s)=r$ as desired. ∎
4 Branching rules
Most of our branching rules can be viewed in terms of guessing that a certain set of vertices $X_{1}$ is in the cover. In the graph $G-X_{1}$, a set of vertices $X_{0}$ may become isolated. We apply rule (P1) to remove $X_{0}$; the resulting subproblem $\langle G-(X_{0}\cup X_{1}),k-|X_{1}|\rangle$ is called principal, and we define the excess to be $|X_{0}|$. Note that the graph $G^{\prime}=G-(X_{0}\cup X_{1})$ may still contain isolated vertices, or have additional simplifications available. We say $\langle G^{\prime},k-|X_{1}|\rangle$ has drop $(\Delta\mu^{\prime},|X_{1}|)$ directly, and the final subproblem $G^{\prime\prime}=\langle G^{\prime\prime},k^{\prime\prime}\rangle$ has drop $(\Delta\mu^{\prime\prime},\Delta k^{\prime\prime})$ after simplification.
Since this plays a central role in our analysis, we define the shadow of a vertex set $X$, denoted $\operatorname{shad}(X)$, to be $\operatorname{minsurp}(G-X)$. For readability, we write $\operatorname{shad}(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for $\operatorname{shad}(\{x_{1},\dots,x_{\ell}\}\cup X_{1}\cup\dots\cup X_{r})=\operatorname{minsurp}(G-\{x_{1},\dots,x_{\ell}\}-(X_{1}\cup\dots\cup X_{r}))$.
Proposition 14.
If $\operatorname{minsurp}(G)\geq 2$, then a principal subproblem $\langle G-X,k^{\prime}\rangle$ with excess $s$ has $\Delta\mu=\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{-}(G^{\prime}))=\tfrac{1}{2}(\Delta k-s+\min\{0,\operatorname{shad}(X)\})$.
Proof.
Here Proposition 4 gives $\lambda(G)=n/2$. Let $G^{\prime}=G-X$, where $|X|=\Delta k+s$. Then $\mu(G^{\prime})=k^{\prime}-\operatorname{\lambda}(G^{\prime})=k^{\prime}-\tfrac{1}{2}(n^{\prime}+\operatorname{minsurp}^{-}(G^{\prime}))$, where $n^{\prime}=n-X=2\lambda(G)-(s+\Delta k)$, which simplifies to $\mu(G^{\prime})=\mu(G)-\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{-}(G^{\prime}))$ as desired. ∎
Observation 15.
Suppose $\operatorname{minsurp}(G)\geq 2$. Then for any indset $I$, there holds $\operatorname{shad}(I)\geq 2-|I|$ and $\operatorname{shad}(N[I])\geq 2-\operatorname{surp}_{G}(I)$.
In particular, for a vertex $u$, there holds $\operatorname{shad}(u)\geq 1$ and $\operatorname{shad}(N[u])\geq 3-\deg(u)$.
Proof.
If $J$ is a min-set of $G-I$, then $2\leq\operatorname{surp}_{G}(J)\leq\operatorname{surp}_{G-I}(J)+|I|$. So $\operatorname{surp}_{G-I}(J)\geq 2-|I|$. Similarly, if $J$ is a min-set of $G-N[I]$, then $2\leq\operatorname{surp}_{G}(I\cup J)=\operatorname{surp}_{G-N[I]}(J)+\operatorname{surp}_{G}(I)$. ∎
By combining Observation 11 and Observation 15, we immediately get the following simplification rule (which is used ubiquitously):
Observation 16.
If $G$ is simplified and $G-\{u,v\}$ has two non-adjacent subquadratic vertices, then $R(u,v)\geq 2$.
Our bread-and-butter branching rule is to choose some vertex $u$, and branch on whether it is in the cover. We refer to this as splitting on $u$. This generates two principal subproblems $\langle G-u,k-1\rangle$ and $\langle G-N[u],k-\deg(u)\rangle$. More generally, in light of Lemma 5, we know that for any critical-set $I$, there is either a good cover containing $I$ or omitting $I$. We can branch on subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-N[I]\rangle$, and we refer to this as splitting on $I$. Splitting on a vertex is a special case where $I$ is a singleton.
Consider the effect of splitting on $u$. By Observation 15, the subproblem $\langle G-u,k-1\rangle$ has drop $(0.5,1)$ and the subproblem $\langle G-N[u],k-\deg(u)\rangle$ has drop $(1,\deg(u))$. The latter bound, however, is very loose and should be improved to get an optimized analysis. In light of Proposition 14, it revolves around the value of $\operatorname{shad}(N[u])$. If $\operatorname{shad}(N[u])\geq 0$, then $\lambda$ drops by $\frac{\deg(u)+1}{2}$ and so the subproblem has drop $(\tfrac{\deg(u)-1}{2},\deg(u))$. This is the “generic” situation for splitting on $u$.
When $\operatorname{shad}(N[u])\leq 0$, we say that $u$ is blocked. If $x$ is a vertex in a min-set of $G-N[u]$, we say that $x$ is a blocker of $u$. Necessarily, $x\nsim u$. This motivates the following powerful branching rule:
(B) If $x$ is a blocker of vertices $u_{1},\dots,u_{\ell}$, then branch on subproblems $\langle G-\{u_{1},\dots,u_{\ell},x\},k-\ell-1\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$.
Proposition 17.
Rule (B) is a valid branching rule.
Proof.
@@ 4 Branching rules
The subproblem $\langle G-N[x],k-\deg(x)\rangle$ is feasible if and only if $G$ has a good cover omitting $x$. The subproblem $\langle G-\{u_{1},\dots,u_{\ell},x\},k-\ell-1\rangle$ is feasible if and only if $G$ has a good cover which includes all vertices $u_{1},\dots,u_{\ell},x$. We claim that, if $G$ is feasible, at least one of these cases holds. For, suppose a good cover of $G$ omits vertex $u_{i}$. Then $G-N[u_{i}]$ has a cover of size at most $k-\deg(u_{i})$. By Lemma 5, this implies that $G-N[u_{i}]$ has such a cover $C^{\prime}$ which omits any min-set $I$ of $G-N[u_{i}]$, and in particular $x\notin C^{\prime}$. Then $G$ has a good cover $C\cup N(u_{i})$ omitting $x$. ∎
@@
In the vast majority of cases where we use (B), we have $\ell=1$, that is, $x$ is a blocker of a single vertex $u$. To provide intuition, keep in mind the picture that if $G-N[u]$ has $s$ isolated vertices $v_{1},\dots,v_{s}$, then $\{v_{1},\dots,v_{s}\}$ would be an indset in $G-N[u]$ with zero neighbors and surplus $-s$, in particular, $\operatorname{shad}(N[u])\leq-s$. These vertices were hidden in the “shadow” of $N[u]$. See Figure 5.
@@
Figure 5: Here, vertices $v_{1},v_{2}$ are both blockers of $u$; in particular, we have $\operatorname{shad}(N[u])\leq-2$.
@@
Also, note that if $\operatorname{shad}(N[u])=0$ exactly, then we could still split on $u$ if desired and get the ideal drop in $\mu$. However, we can also apply rule (B) if desired, and this is usually more profitable. Vertices with $\operatorname{shad}(N[u])=0$ share many properties with vertices with $\operatorname{shad}(N[u])<0$ strictly. This is the reason for the somewhat unintuitive definition of blocked vertex.
5 Analysis of blockers and splitting
@@
We will develop a series of branching rules (typically using rule (B)) to handle blocked vertices, along with a number of other related cases. The most significant consequence of these rules is that we are able to get near-ideal branch sequences for high-degree vertices. Specifically, we will show the following main result which is the heart of our branching algorithm:
@@
theoremthmA Suppose $G$ is simplified, and let $r=\operatorname{maxdeg}(G)$. Depending on $r$, then following branch-seqs are available:
@@
The proof of Section 5 has many cases which will require us to build up slowly from analysis of low-degree vertices. For Section 5 only, we always assume that the starting graph $\boldsymbol{G}$ is simplified. We begin with a few elementary observations.
Observation 18.
Suppose $u$ is blocked and let $I$ be a min-set of $G-N[u]$. There is at least one vertex $x\in I$ with $\operatorname{codeg}(x,u)\geq 1$.
@@
If not, we would have $\operatorname{surp}_{G}(I)=\operatorname{surp}_{G-N[u]}(I)\leq 0$, contradicting that $G$ is simplified. ∎
@@
Suppose vertex $u$ has $\operatorname{shad}(N[u])\leq 4-\deg(u)$ and $x$ is a blocker of $u$. If we apply (B) to vertices $u,x$, then subproblem $G^{\prime}=\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$. Moreover, if $\deg(x)=3$ and $x$ is contained in a non-singleton min-set of $G-N[u]$, then $G^{\prime}$ has drop $(1,5)$.
@@
Since $G$ is simplified, we have $\deg(x)\geq 3$ and $\operatorname{minsurp}(G)\geq 2$. If $\deg(x)=4$, then $G^{\prime}$ has drop $(1,4)$ directly. So, suppose $\deg(x)=3$ exactly. Consider any min-set $I$ of $G-N[u]$ with $x\in I$. The indset $J=I\cup\{u\}\setminus\{x\}$ has $\operatorname{surp}_{G-N[x]}(J)\leq\operatorname{surp}_{G-N[u]}(I)+(\deg(u)-\deg(x))\leq 1$. On the other hand, $\operatorname{shad}(N[x])\geq\operatorname{minsurp}(G)-\deg(x)+1\geq 0$. So Lemma 13 gives $R(N[x])\geq|J|=|I|$.
By Proposition 14 we see that $G^{\prime}$ has drop $(1,3)$ directly. Since $|I|\geq 1$ trivially, we have $R(N[x])\geq 1$ and $G^{\prime}$ has drop $(1,4)$ after simplification. If $|I|\geq 2$, then $R(N[x])\geq 2$ and likewise $G^{\prime}$ has drop $(1,5)$ after simplification. ∎
5.1 Branching rules for indsets with surplus two.
It will be very important to develop special branching rules for indsets with surplus two. Since $G$ is assumed to be simplified, note that any such set $I$ will be a critical set, and we can split on $I$ to get subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-|I|-2\rangle$. However, this branching rule is not powerful enough for us; we develop specialized branching rules for smaller indsets (indsets with cardinality two or three).
Proposition 20.
Suppose $G$ has an indset $I$ with $\operatorname{surp}_{G}(I)=2$, and let $z\in N(I)$. If we split on $z$, the subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,1+|I|)$. If $z$ has a blocker $x\notin I$ and we apply (B) to $z,x$, the subproblem $\langle G-\{z,x\},k-2\rangle$ has drop $(1,2+|I|)$.
Proof.
Subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,1)$ directly. We have $\operatorname{surp}_{G-z}(I)=\operatorname{surp}_{G}(I)-1=1$, and $\operatorname{shad}(G-z)\geq\operatorname{minsurp}(G)-1\geq 1$. So (P2) applied to $I$ gives $R(z)\geq|I|$. Likewise, subproblem $\langle G-\{z,x\},k-2\rangle$ has drop $(1,2)$ directly, and $\operatorname{shad}(z,x)\geq\operatorname{minsurp}(G)-2\geq 0$ and $\operatorname{surp}_{G-\{z,x\}}(I)\leq\operatorname{surp}_{G}(I)-1\leq 1$. So Lemma 13 gives $R(z,x)\geq|I|$. ∎
@@
If $|I|\geq 4$, then splitting on $I$ gives branch-seq $[(1,4),(1,6)]$. So suppose $|I|=3$. We first claim that there is some vertex $z$ with $|N(z)\cap I|=2$. For, consider the set of vertices $J=\{z:|N(z)\cap I|\geq 2\}$. The vertices in $I$ are independent and have degree at least three, so there are at least 9 edges from $I$ to $N(I)$. Since $|N(I)|=5$, by the pigeonhole principle we must have $|J|\geq 2$. If $N(z)\not\subseteq I$ for all $z\in J$, then $J$ would be an indset with surplus at most one, contradicting that $G$ is simplified. Hence there is $z\in J$ with $|N(z)\cap I|=2$ exactly.
Let $z$ be any such vertex, and let $I=\{x_{1},x_{2},y\}$ where $z\sim x_{1},z\sim x_{2},z\not\sim y$ and $\deg(x_{1})\leq\deg(x_{2})$. If $\deg(z)\leq 4$, then we split on $I$, generating subproblems $\langle G-I,k-3\rangle$ and $\langle G-N[I],k-5\rangle$ with drops $(1,3)$ and $(1,5)$ directly, where $G-I$ has a subquadratic vertex $z$ so it has drop $(1,4)$ after simplification. If $\deg(z)\geq 5$ and $\operatorname{shad}(N[z])\geq 5-\deg(z)$, then we split on $z$; by Proposition 20, subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,4)$, and subproblem $\langle G-N[z],k-\deg(z)\rangle$ has drop $(2,5)$ directly. If $\deg(z)\geq 5$ and $z$ has a blocker $t\neq y$, then we apply (B) to $z,t$; by Proposition 19, subproblem $\langle G-N[t],k-\deg(t)\rangle$ has drop $(1,4)$ and by Proposition 20, subproblem $\langle G-\{z,t\},k-2\rangle$ has drop $(1,5)$.
The only remaining possibility is if $\deg(z)\geq 5$ and $y$ is the only blocker of $z$, i.e. $\deg(z)=5,\deg(y)=3$ and $N(y)\subseteq N(z)$. Each vertex $x_{i}$ has $\deg(x_{i})+\deg(y)-\operatorname{codeg}(x_{i},y)\leq|N(I)|$, i.e. $\operatorname{codeg}(x_{i},y)\geq\deg(x_{i})-2$. If $\deg(x_{1})=3$, this implies that $x_{1},y$ share some neighbor $u$; but in this case $G$ has a 3-triangle $x_{1},u,z$, contradicting that $G$ is simplified.
So we suppose $\deg(x_{2})\geq\deg(x_{1})\geq 4$ and then $\operatorname{codeg}(x_{1},y)\geq 2$ and $\operatorname{codeg}(x_{2},y)\geq 2$. Equivalently, $\deg_{G-N[x_{1}]}(y)\leq 1$ and $\deg_{G-N[x_{2}]}(y)\leq 1$. So $y$ is a blocker of all three vertices $x_{1},x_{2},z$. We apply (B) giving subproblems $\langle G-N[y],k-\deg(y)\rangle$ and $\langle G-\{x_{1},x_{2},z,y\},k-4\rangle$, with drops $(1,3)$ and $(1,4)$ directly. Furthermore, $\operatorname{surp}_{G-N[y]}(\{x_{1},x_{2}\})\leq 2-|N_{G-N[y]}(I)|=2-2=0$ and neither $x_{1}$ or $x_{2}$ is isolated in $G-N[y]$ (they retain neighbor $z$). So we can apply (P1) to indset $\{x_{1},x_{2}\}$ in $G-N[y]$, giving $R(N[y])\geq 2$. So $G-N[y]$ has drop $(1,5)$ after simplification. ∎

### v7

@@ 2.2 Branching framework
We say branch-seq $B^{\prime}$ dominates $B$ if $\operatorname{val}_{a,b}(B^{\prime})\leq\operatorname{val}_{a,b}(B)$ for all $a,b\geq 0$, and $G$ has a branch-seq $B$ available if we can efficiently find a branching rule with a branch-seq dominating $B$.
Most of our branching rules come from guessing that a certain vertex set $X$ is in the cover, and then removing a set of vertices $Y$ which are isolated in GXG-X. The resulting subproblem $\langle G^{\prime},k^{\prime}\rangle$ with $G^{\prime}=G-(X\cup Y),k^{\prime}=k-|X|$ is called principal with $\Delta k=|X|$ and excess $s=|Y|$. The graph $G^{\prime}$ may have other simplifications available. We say $\langle G^{\prime},k^{\prime}\rangle$ has drop $(\Delta\mu^{\prime},|X|)$ directly, and the final subproblem $G^{\prime\prime}=\langle G^{\prime\prime},k^{\prime\prime}\rangle$ has drop $(\Delta\mu^{\prime\prime},\Delta k^{\prime\prime})$ after simplification. Note that if we apply (P1) to $G^{\prime}$, then the subproblem $G^{\prime\prime}=\langle G-(X\cup Y\cup N[Z]),k-|X|-|N(Z)|\rangle$ is again principal, with excess $|Y|+|Z|$; however, if we apply (P2) or (P3) to $G^{\prime}$, then $G^{\prime\prime}$ is not principal as it requires other operations such as adding edges.
Proposition 16.
If $\operatorname{minsurp}(G)\geq 0$, then a principal subproblem $\langle G-Z,k-\Delta k\rangle$ with excess $s$ has $\Delta\mu=\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{0}(G-X))$.
@@
Proposition 4 gives $\lambda(G)=n/2$. Let $G^{\prime}=G-Z$, where $|Z|=\Delta k+s$. Then $\mu(G^{\prime})=k^{\prime}-\operatorname{\lambda}(G^{\prime})=k^{\prime}-\tfrac{1}{2}(n^{\prime}+\operatorname{minsurp}^{0}(G^{\prime}))$, where $n^{\prime}=n-|Z|=2\lambda(G)-(s+\Delta k)$, which simplifies to $\mu(G^{\prime})=\mu(G)-\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{0}(G^{\prime}))$. ∎
Consider a critical-set II of $G$. By Lemma 6, we know that there is either a good cover containing $I$, or disjoint to $I$. In the latter case, all the neighbors of $I$ must go into the cover. There is a natural branching rule to guess which of these possibilities holds:
(B-Crit): For a critical-set $I$ branch on subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-|N(I)|\rangle$.
@@
Suppose that $\operatorname{minsurp}(G)\geq 2$. Let $I$ be a critical-set with $\operatorname{surp}_{G}(I)=\ell,|I|=r$. Then (B-Crit) generates principal subproblems $G_{1}=\langle G-I,k-r\rangle$ and $G_{0}=\langle G-N[I],k-r$-$\ell\rangle$ with excess 0 and $r$ respectively. Furthermore:
  • If $r=1$, then the subproblem $G_{1}$ has $\Delta\mu=\tfrac{1}{2}$.
  • If $r>1$, then the subproblem $G_{1}$ has $\Delta\mu=\tfrac{1}{2}(r+\operatorname{minsurp}^{0}(G-I))\geq 1$.
  • The subproblem $G_{0}$ has $\Delta\mu=\tfrac{1}{2}(\ell+\operatorname{minsurp}^{0}(G-N[I]))\geq 1$.
@@
The bound $\Delta\mu=\tfrac{1}{2}(r+\operatorname{minsurp}^{0}(G-I))$ for $G_{1}$ follows from Proposition 16. Also, by Observation 9, we have $\operatorname{minsurp}(G-I)\geq\operatorname{minsurp}(G)-|I|\geq 2-r$. Thus, if $r=1$, we have $\operatorname{minsurp}^{0}(G-I)=0$. Similarly, by Proposition 16, we have $\Delta\mu=\tfrac{1}{2}(\ell+\operatorname{minsurp}^{0}(G-N[I]))$ and by Observation 9, we have $\operatorname{minsurp}(G-N[I])\geq 2-\operatorname{surp}_{G}(I)=2$-$\ell$. ∎
3 Analysis of vertex branching
Our bread-and-butter branching rule is to choose a vertex $u$ and branch on whether it is in the cover. We refer to this as splitting on $u$. It is a special case of (B-Crit) with a singleton set $I=\{u\}$, generating principal subproblems $\langle G-u,k-1\rangle$ and $\langle G-N[u],k-\deg(u)\rangle$. By Proposition 16, if we start with a simplified graph, then the subproblems have drops $(0.5,1)$ and $(1,\deg(u))$ respectively.
Let us examine the subproblem $\langle G-N[u],k-\deg(u)\rangle$ more carefully: if $\operatorname{shad}(N[u])\geq 0$, then the subproblem has drop $(\tfrac{\deg(u)-1}{2},\deg(u))$. This is the “generic” and desired situation. Otherwise, when $\operatorname{shad}(N[u])\leq 0$, we say that $u$ is blocked. If $x$ is a vertex in a min-set of GN[u]G-N[u], we say that $x$ is a blocker of $u$. This motivates the following powerful branching rule:
(B-Block): Given a vertex $x$ which is a blocker of each of the vertices $u_{1},\dots,u_{\ell}$, branch on subproblems $G_{1}=\langle G-\{u_{1},\dots,u_{\ell},x\},k-\ell-1\rangle$ and $G_{0}=\langle G-N[x],k-\deg(x)\rangle$.
Proposition 18.
(B-Block) is a valid branching rule.
@@ 3 Analysis of vertex branching
The subproblem $G_{0}$ is feasible if and only if $G$ has a good cover omitting $x$, and the subproblem $G_{1}$ is feasible if and only if $G$ has a good cover which includes all vertices $u_{1},\dots,u_{\ell},x$. We claim that, if $\langle G,k\rangle$ is feasible, at least one of these cases holds. For, suppose a good cover of $G$ omits vertex $u_{i}$. Then $G-N[u_{i}]$ has a cover of size at most $k-\deg(u_{i})$. By Lemma 6, then $G-N[u_{i}]$ has such a cover $C^{\prime}$ which omits any min-set II of $G-N[u_{i}]$, and in particular $x\notin C^{\prime}$. Then $G$ has a good cover $C=C^{\prime}\cup N(u_{i})$ omitting $x$. ∎
@@
In the vast majority of cases, we use (B-Block) with $\ell=1$, that is, $x$ is a blocker of a single vertex $u$. For intuition, keep in mind the picture where GN[u]G-N[u] has $s$ isolated vertices $v_{1},\dots,v_{s}$, so that $\{v_{1},\dots,v_{s}\}$ is an indset in GN[u]G-N[u] with zero neighbors and surplus s-s, in particular, $\operatorname{shad}(N[u])\leq-s$. See Figure 6.
@@
Figure 6: Here, vertices $v_{1},v_{2}$ are both blockers of $u$; in particular, $\operatorname{shad}(N[u])\leq-2$.
@@
Also, note that if $\operatorname{shad}(N[u])=0$ exactly, then we could still split on $u$ if desired and get the ideal drop in $\mu$. However, we can also apply rule (B-Block), and this is usually more profitable. Vertices with $\operatorname{shad}(N[u])=0$ share many properties with vertices with $\operatorname{shad}(N[u])<0$ strictly. This is the reason for the somewhat unintuitive definition of blocked vertex.
@@
We will develop a series of branching rules to handle blocked vertices and other related cases. The most significant consequence of these rules is to give near-ideal branch sequences for high-degree vertices. Specifically, we have the following main result which is the heart of our algorithm:
@@
theoremthmA Suppose $G$ is simplified, and let $r=\operatorname{maxdeg}(G)$. Depending on $r$, the following branch-seqs are available:
@@
The proof of Figure 6 has many cases which we will build up slowly. For Section 3 only, we always assume that the starting graph $\boldsymbol{G}$ is simplified.
3.1 Branching rules for indsets with surplus two
We first develop branching rules for non-singleton indsets with surplus two. As a starting point, note that since $G$ is assumed to be simplified, we have $\operatorname{minsurp}(G)\geq 2$ hence such a set $I$ would be a critical-set. We can apply (B-Crit) to obtain subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-r-2\rangle$. These are principal subproblems, and they have drops $(1,r),(1,r+2)$ respectively.
This is already a powerful branching rule, but it can be improved when $|I|=2$ or $|I|=3$ by finding additional simplifications in the subproblems. We start with a few preliminary observations.
Observation 19.
Suppose $G$ has an indset $I$ with $\operatorname{surp}_{G}(I)=2$, and let $z\in N(I)$.
Then $S(z)\geq|I|$, and for any vertex $x\notin I$, there holds $S(z,x)\geq|I|$.
@@ 3.1 Branching rules for indsets with surplus two
We have $\operatorname{surp}_{G-z}(I)=\operatorname{surp}_{G}(I)-1=1$, and $\operatorname{shad}(z)\geq\operatorname{minsurp}(G)-1\geq 1$. So (P2) applied to $I$ gives $S(z)\geq|I|$. Likewise, $\operatorname{shad}(z,x)\geq\operatorname{minsurp}(G)-2\geq 0$ and $\operatorname{surp}_{G-\{z,x\}}(I)\leq\operatorname{surp}_{G}(I)-1\leq 1$. So Lemma 13 gives $S(z,x)\geq|I|$. ∎
@@
Suppose $\operatorname{shad}(N[u])\leq 4-\deg(u)$ and $x$ is a blocker of $u$. If we apply (B-Block) to vertices $u,x$, then subproblem $G^{\prime}=\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$. Moreover, if $\deg(x)=3$ and $x$ is contained in a non-singleton min-set of GN[u]G-N[u], then $G^{\prime}$ has drop $(1,5)$.
@@
If $\deg(x)=4$, then $G^{\prime}$ has drop $(1,4)$ directly. So, suppose $\deg(x)=3$; by Proposition 16, $G^{\prime}$ has drop $(1,3)$ directly. Let $I\ni x$ be a min-set of GN[u]G-N[u]. The indset $J=I\cup\{u\}\setminus\{x\}$ has $\operatorname{surp}_{G-N[x]}(J)\leq\operatorname{surp}_{G-N[u]}(I)+(\deg(u)-\deg(x))\leq 1$. On the other hand, $\operatorname{shad}(N[x])\geq\operatorname{minsurp}(G)-\deg(x)+1\geq 0$. So Lemma 13 gives $S(N[x])\geq|J|=|I|$. ∎
@@
If $|I|\geq 4$, then we simply apply (B-Crit). So suppose $|I|=3$. The vertices in $I$ are independent, so there are at least $\sum_{x\in I}\deg(x)\geq 9$ edges from $I$ to $N(I)$. Since $|N(I)|=5$, by the pigeonhole principle there must be at least one vertex $z$ with $|N(z)\cap I|\geq 2$. Our strategy will revolve around a chosen vertex $z$; as a tie-breaking rule, we assume that in a given case analysis there were no other vertices in a previous case.
Case I: $\boldsymbol{\deg(z)\leq 4}$ and $\boldsymbol{N(z)\not\subseteq I}$. We apply (B-Crit) to $I$, generating subproblems $\langle G-I,k-3\rangle$ and $\langle G-N[I],k-5\rangle$ with drops $(1,3)$ and $(1,5)$ directly. The former subproblem has a subquadratic vertex $z$ so it has drop $(1,4)$ after simplification.
Case II: $\boldsymbol{\deg(z)\geq 5}$ and $\boldsymbol{\operatorname{shad}(N[z])\geq 5-\deg(z)}$. We split on $z$; by Observation 19, subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,4)$ after simplification, and subproblem $\langle G-N[z],k-\deg(z)\rangle$ has drop $(2,5)$ directly.
Case III: $\boldsymbol{\deg(z)\geq 5}$ and $\boldsymbol{z}$ has a blocker $\boldsymbol{t\notin I}$. We apply (B-Block) to $z,t$; by Proposition 20, subproblem $\langle G-N[t],k-\deg(t)\rangle$ has drop $(1,4)$ and by Observation 19, subproblem $\langle G-\{z,t\},k-2\rangle$ has drop $(1,5)$ after simplification.
Case IV: $\boldsymbol{\deg(z)\geq 5}$. Let $I=\{x_{1},x_{2},y\}$ where $z\sim x_{1},z\sim x_{2}$. Since we are not in Case II or Case III, it must be that $y$ is the sole blocker of $z$: that is, we have $N(y)\subseteq N(z)$.
Each vertex $x_{i}$ has $|N(x_{i},y)|=\deg(x_{i})+\deg(y)-\operatorname{codeg}(x_{i},y)\leq|N(I)|$, i.e. $\operatorname{codeg}(x_{i},y)\geq\deg(x_{i})-2\geq 1$. So we must have $\deg(x_{1}),\deg(x_{2})\geq 4$, as otherwise some $x_{i}$ would have a 3-triangle with $u,z$, contradicting that $G$ is simplified. As a result, there are at least $11$ edges going from $I$ to $N(I)$; since $|N(I)|=5$ the pigeonhole principle implies there must be some vertex $z^{\prime}$ with $|N(z^{\prime})\cap I|=3$. Since $z^{\prime}\in N(y)\subseteq N(z)$, we must have $z^{\prime}\sim z$ as well. So $N(z^{\prime})\not\subseteq I$. Necessarily, any blocker of this vertex $z^{\prime}$ would be outside $I$. Depending on the degree of $z^{\prime}$, it would fall into one of the earlier cases. This contradicts our tie-breaking rule for $z$. So Case IV is impossible.
Figure 7: Illustration of Case IV
Case V: $\boldsymbol{N(z)\subseteq I}$. Necessarily $\deg(z)=3$. There are at least $6$ edges from $I$ to $N(I)\setminus\{z\}$, so by the pigeonhole principle there is another vertex $z^{\prime}\neq z$ with $|N(z^{\prime})\cap I|\geq 2$. If $N(z^{\prime})\subseteq I$, then necessarily $z\not\sim z^{\prime}$. So $\{z,z^{\prime}\}$ is an indset of surplus one, contradicting that $G$ is simplified. If $N(z^{\prime})\not\subseteq I$, it would fall into one of the earlier cases. In either event, Case V is also impossible. ∎